### v1

# Lesser Green’s Function and Chirality Entanglement Entropy via the In-Medium NJL Model

Abstract
We study chiral symmetry restoration in hot-and-dense quark matter using the von Neumann chirality entropy within the in-medium Nambu-Jona-Lasinio (NJL) model. Starting from the lesser Green’s function $G^{<}(k)$, the reduced correlator $C_{L}=P_{L}G^{<}P_{L}$ is constructed, and the associated entropy $S_{\chi}=-\mathrm{Tr}[C_{L}\ln C_{L}+(1-C_{L})\ln(1-C_{L})]$ is introduced to quantify the entanglement between left- and right-handed quark sectors. The dynamical quark mass $M_{q}(T,\mu_{q})$ obtained from the gap equation exhibits the expected QCD-like phase structure: A second-order transition in the chiral limit and a smooth crossover for finite $m_{q}$. The chirality entropy $S_{\chi}$ increases monotonically with temperature and chemical potential, approaching a maximal value as $M_{q}\to 0$. We also explore the critical exponents and scaling behavior of those quantities, yielding $\beta_{S_{\chi}}\simeq 1$. This demonstrates that $S_{\chi}$ serves as an information-theoretic probe for chiral symmetry restoration, linking dynamical mass generation to quantum entanglement in strongly interacting matter.
Chiral symmetry restoration, Nambu–Jona-Lasinio model, von Neumann entropy, chirality entanglement, lesser Green’s function, critical exponent and scaling behavior, quantum information in QCD
preprint: PKNU-NuHaTh-2025
I Introduction
The spontaneous breakdown (SBCS) and restoration of chiral symmetry constitute a fundamental aspect of Quantum Chromodynamics (QCD) and play a central role in the mass generation mechanism of hadrons and in the formation of the quark-gluon plasma (QGP) at high temperature or density [1, 2, 3, 4, 5]. In the QCD vacuum, chiral symmetry is dynamically broken, leading to a large constituent quark mass and giving rise to light pseudoscalar mesons as Nambu-Goldstone bosons. As temperature or baryon density increases, the quark condensate $\langle\bar{q}q\rangle$ melts, the constituent mass decreases, and chiral symmetry is partially restored. Lattice QCD simulations have established that this restoration occurs through a smooth crossover at vanishing chemical potential for physical quark masses [6, 7]. At the same time, effective models such as the Nambu–Jona-Lasinio (NJL), Polyakov-NJL (PNJL), and instanton-based models successfully reproduce the qualitative features of the QCD phase diagram [2, 3, 8, 9, 10].
Recently, concepts from quantum information theory have been applied to strongly interacting systems to characterize their microscopic coherence and correlations in terms of entanglement entropy [11, 12, 13, 14]. In QCD and other field-theoretical systems, entanglement-based observables have been used to probe confinement, deconfinement, topological ordering, and spin structures [15, 16, 17, 18]. The von Neumann entropy, in particular, provides a quantitative measure of quantum coherence loss and has been proposed as a diagnostic tool for phase transitions in gauge theories and many-body systems.
In this work, we formulate the von Neumann chirality entropy as an information-theoretic probe of chiral symmetry restoration within the in-medium NJL model. Starting from the lesser Green’s function $G^{<}(k)$ in the real-time formalism [19, 20], we construct the reduced left-handed correlator $C_{L}(\bm{k})=P_{L}G^{<}(k)P_{L}$ and define the entropy
$S_{\chi}=-\mathrm{Tr}\,[C_{L}\ln C_{L}+(1-C_{L})\ln(1-C_{L})].$
This entropy measures the degree of quantum entanglement between left- and right-handed quark sectors, which is dynamically generated through the mass term $M_{q}(\bar{q}_{L}q_{R}+\bar{q}_{R}q_{L})$. In the chirally broken phase ($M_{q}\neq 0$), the left and right components are polarized and coherence is maintained, while in the restored phase ($M_{q}\to 0$), they become maximally mixed, leading to a large $S_{\chi}$.
We demonstrate that $S_{\chi}(T,\mu_{q})$ increases monotonically with both temperature and chemical potential, reflecting the progressive loss of chiral coherence and the transition to a maximally mixed state. The resulting behavior mirrors the thermodynamic phase structure of the NJL model: A second-order transition in the chiral limit and a smooth crossover for finite current quark mass. Thus, the von Neumann chirality entropy serves as an information-theoretic probe that connects the dynamical mechanism of mass generation to the microscopic structure of quantum entanglement in QCD matter.
Most existing entanglement-related studies in QCD address spatial bipartition entropy, Rényi entropies, or color entanglement in Euclidean or lattice formulations. In contrast, the present work develops an information-theoretic quantity associated with the chiral internal degree of freedom, constructed directly from the real-time lesser Green function. The resulting chirality-reduced entropy captures dynamical changes in the quasiparticle spectrum and offers a new, complementary perspective on the chiral phase transition beyond conventional order parameters and previously studied entanglement measures.
The remainder of this paper is organized as follows. In Sec. II, we outline the theoretical framework of the finite-temperature Nambu-Jona-Lasinio model and introduce the lesser Green’s function in the real-time formalism, which provides the foundation for the subsequent definition of the reduced correlator. In Sec. III, we derive the left-handed reduced correlator $C_{L}(\bm{k})$ and analyze its connection to the dynamical mass and chiral mixing. In Sec. IV, the von Neumann chirality entropy is formulated and its physical interpretation as an information-theoretic measure of chiral entanglement is discussed in detail. Section V presents the numerical results for the dynamical quark mass and chirality entropy as functions of temperature and chemical potential, together with a detailed physical discussion of the QCD-like phase structure. Finally, Sec. VI summarizes the main conclusions and remarks on possible extensions of the present approach, including applications to the PNJL model and to lattice-QCD analyses of the entanglement in hadronic matter.
II In-medium NJL Model
The NJL model is based on a four-fermion interaction that preserves global chiral symmetry in the massless limit. Its Lagrangian reads:
$\mathcal{L}_{\text{NJL}}=\bar{q}(i\gamma^{\mu}\partial_{\mu}-m_{q})q+G\left[(\bar{q}q)^{2}+(\bar{q}i\gamma_{5}\bm{\tau}q)^{2}\right],$ (1)
where $q=(u,d)^{T}$ represents the light-flavor quark fields, $m_{q}$ is the current quark mass (assumed to be small but nonzero), and $G$ is a dimensionful coupling constant characterizing the strength of the effective scalar and pseudoscalar interactions. The second term in the Lagrangian contains chirally invariant combinations of bilinear quark operators, reflecting the underlying $SU(2)_{L}\times SU(2)_{R}$ symmetry.
The NJL model does not confine quarks, yet it successfully captures dynamical mass generation. When the interaction strength $G$ exceeds a critical value, the vacuum becomes unstable against the formation of a quark condensate $\langle\bar{q}q\rangle$, which serves as an order parameter for SBCS. This results in the generation of a constituent quark mass $M_{q}$, given by the gap equation:
$M_{q}=m_{q}-2G\langle\bar{q}q\rangle.$ (2)
Here, the dynamical mass $M_{q}$ significantly exceeds the current mass $m_{q}$, reflecting the nonperturbative nature of the QCD vacuum. This mechanism also leads to the emergence of massless Nambu-Goldstone bosons, identified with the pions in the chiral limit, associated with the broken axial generators of the chiral symmetry. The NJL model thus provides a minimal yet powerful framework for exploring the interplay between chiral symmetry and the fermionic vacuum structure in low-energy QCD. The conventional NJL model parameters are employed in the numerical calculations and given in Table 1.
$m_{q}$ [MeV] $\Lambda$ [MeV] $G\Lambda^{2}$
5.25 631.4 2.14
Table 1: NJL parameter set used in the present study.
Within the Matsubara formalism, this temperature-dependent $M_{q}(T,\mu)$ at finite quark chemical potential $\mu_{q}$ is evaluated as
$M_{q}(T,\mu)=m_{q}-2G\langle\bar{q}{q}\rangle,\,\,\,\langle\bar{q}{q}\rangle=-4N_{c}N_{f}\int^{\Lambda}_{0}\frac{d^{3}\vec{k}}{(2\pi)^{3}}\frac{M_{q}}{E_{k}}\big[1-n_{+}-n_{-}\big],$ (3)
where $E^{2}_{k}=|\vec{k}|^{2}+M_{q}^{2}$, and the Fermi-Dirac distribution function reads $n_{F}(E_{k}\pm\mu)=1/[e^{(E_{k}\pm\mu)/T}+1]\equiv n_{\mp}$. As $T$ increases, thermal fluctuations favor quark-antiquark pair excitations and diminish the condensate magnitude, thereby reducing the dynamically generated mass $M_{q}$ and softening the symmetry-breaking scale, signaling the partial restoration of chiral symmetry in terms of the crossover for finite $m_{q}$. At the same time, the second-order phase transition occurs in the chiral limit $m_{q}=0$.
Fig. 1 shows the temperature-chemical potential $(T,\mu_{q})$ dependence of the dynamical quark mass $M_{q}(T,\mu_{q})$ obtained from the NJL gap equation, Eq. (3). The color map shows the magnitude of $M_{q}$, which indicates the degree of spontaneous chiral symmetry breaking. Fig. 1(a) corresponds to the chiral limit ($m_{q}=0$), while Fig. 1(b) shows the case with a small explicit current quark mass ($m_{q}=5.25~\mathrm{MeV}$).
$M_{q}$ takes a considerable value ($M_{q}\simeq 300$$350~\mathrm{MeV}$) near $T=\mu_{q}=0$, indicating that chiral symmetry is strongly broken in the vacuum. As $T$ increases at $\mu_{q}=0$, the constituent mass decreases continuously to zero, signaling a second-order phase transition to the chirally symmetric phase. Along the $\mu_{q}$ axis, the same qualitative behavior occurs: $M_{q}$ decreases with increasing $\mu_{q}$, but the transition sharpens and becomes first order at high densities. The region where the continuous and discontinuous transitions meet corresponds to the critical end point (CEP), a hallmark of the QCD-like phase structure reproduced in the NJL framework. Thus, Fig. 1(a) encapsulates the essential thermodynamic mechanism of chiral symmetry restoration through the temperature- and density-induced melting of the quark condensate as in Eq. (2).
In Fig. 1(b), the inclusion of a finite $m_{q}$ explicitly breaks chiral symmetry, transforming the sharp phase boundary of the chiral limit into a smooth crossover. Because the explicit breaking term prevents $M_{q}$ from vanishing entirely, the transition is gradual rather than singular: $M_{q}$ decreases steadily with increasing $T$ and $\mu_{q}$, remaining finite ($M_{q}\simeq 100$ - $150~\mathrm{MeV}$) even in the high-temperature region. This continuous crossover behavior is consistent with lattice QCD results at physical quark masses, confirming that the NJL model correctly captures the qualitative pattern of chiral symmetry restoration in hot-and-dense matter.
            Refer to caption  (a)                         Refer to caption  (b)
Figure 1: (a) QCD phase diagram of the NJL model, denoted by the effective quark mass $M_{q}$, as a function of $T$ and $\mu_{q}$ for $m_{q}=0$ (a) and $m_{q}=5.25$ MeV (b).
A variation of the model parameters within reasonable ranges, $\Lambda=(630\pm 30)\,\mathrm{MeV}$ and $G\Lambda^{2}=2.1\pm 0.1$, was also tested and found to change only the overall magnitude of the numerical results slightly without altering its qualitative behavior or the position of the pseudo-critical region. Hence, the physical trends discussed above are robust against moderate parameter variations.
III Lesser Green’s Function and Chiral Correlator Framework
In the Schwinger-Keldysh real-time formalism [20], two correlation functions are defined as
$G^{>}(x,y)=-i\langle q(x)\bar{q}(y)\rangle,\,\,\,\,G^{<}(x,y)=i\langle\bar{q}(y)q(x)\rangle.$ (4)
The symbols >> and << denote contour ordering along the Keldysh time path. Physically, $G^{<}$ measures correlations among occupied states and serves as the momentum-space one-body density matrix.
In equilibrium, these are related to the spectral function $A(p)$ and the Fermi distribution $n_{F}(p^{0})$:
$G^{<}(p)=i\,n_{F}(p^{0})A(p),\,\,\,\,G^{>}(p)=-i[1-n_{F}(p^{0})]A(p),$ (5)
where $A(p)=i[G^{R}(p)-G^{A}(p)]$. Thus, $G^{<}$ directly encodes the occupied part of the spectral density.
Within the mean-field NJL model, the lesser propagator is
$G^{<}(k)=2\pi i\Big[n_{+}\,\Lambda_{+}(\bm{k})\,\delta(k_{0}-E_{k})+(1-n_{-})\,\Lambda_{-}(\bm{k})\,\delta(k_{0}+E_{k})\Big],$ (6)
The positive $(+)$ and negative ()(-) energy projectors are given by
$\Lambda_{\pm}(\bm{k})=\frac{1}{2}\left[1\pm\frac{\gamma^{0}(\bm{\gamma}\cdot\bm{k}+M_{q})}{E_{k}}\right].$ (7)
They satisfy $\Lambda_{\pm}^{2}=\Lambda_{\pm}$, $\Lambda_{+}\Lambda_{-}=0$, $\Lambda_{+}+\Lambda_{-}=1$, and
[…]
using ther energy eigenvalue euqations $\gamma^{0}u_{\pm}(\bm{k})=\pm\,u_{\pm}(\bm{k})$. Integrating over $k_{0}=\pm E_{k}$ gives a compact on-shell representation:
$G^{<}(k)=n_{+}(E_{k})\,\Lambda_{+}(\bm{k})+[1-n_{-}(E_{k})]\,\Lambda_{-}(\bm{k}).$ (9)
[…]
IV Chirality Decomposition and Reduced Correlator
The Dirac field can be decomposed into its left- and right-handed components,
$q=q_{L}+q_{R},\,\,\,\,q_{L,R}=P_{L,R}q,\,\,\,\,P_{L,R}=\frac{1\mp\gamma^{5}}{2}.$ (10)
Here, $P_{L,R}$ are projection operators in chiral space that separate fermionic degrees of freedom with definite handedness. The chiral symmetry of the Lagrangian ensures that, in the absence of a mass term, the left- and right-handed fields evolve independently. The mass term introduces mass-induced mixing between left- and right-handed components of the Dirac field, so that a nonzero constituent mass $M_{q}$ acts as a quantum mixing source linking the two sectors. In the chiral limit $M_{q}\to 0$, the mass coupling vanishes, and the Lagrangian decouples into independent left- and right-handed parts, leading to no direct mass-induced entanglement.
To quantify this coupling, we construct the left-handed reduced correlator by projecting the lesser Green’s function $G^{<}(k)$ onto the left-handed subspace:
$C_{L}(\bm{k})=P_{L}G^{<}(k)P_{L}=n_{+}\,P_{L}\Lambda_{+}(\bm{k})P_{L}+(1-n_{-})\,P_{L}\Lambda_{-}(\bm{k})P_{L}.$ (11)
Here $n_{\pm}=n_{F}(E_{k}\mp\mu_{q})$ are the Fermi-Dirac distribution functions, and $\Lambda_{\pm}(\bm{k})$ denote the positive- and negative-energy projectors as discussed previously. The first term represents the contribution of occupied quark states, while the second encodes the vacuum and antiquark components. Thus $C_{L}(\bm{k})$ contains information on how many left-handed states are occupied and how strongly they are mixed with right-handed ones through the mass term. $C_{L}$ is used as the projected correlation matrix, and it may clarify that it is not the thermal Bloch matrix but the Gaussian reduced state. Justification for using $C_{L}=P_{L}G^{<}P_{L}$ as a reduced density matrix is discussed in Appendix B.
            Refer to caption  (a)                         Refer to caption  (b)
Figure 2: Effective occupation probability for $q_{L}$, $\nu_{k}$ in Eq. (13) for $M_{q}=0.1$ GeV (a) and $0.3$ GeV (b) with $|\bm{k}|=1$ GeV.
Using Eq. (7) in the Weyl basis yields
$P_{L}\Lambda_{+}(\bm{k})P_{L}=\frac{1}{2}\left(1-\frac{M_{q}}{E_{k}}\right)P_{L},\,\,\,\,P_{L}\Lambda_{-}(\bm{k})P_{L}=\frac{1}{2}\left(1+\frac{M_{q}}{E_{k}}\right)P_{L}.$ (12)
These relations demonstrate that the mass term $M_{q}$ controls the weight of left-handed components in positive- and negative-energy sectors. When $M_{q}/E_{k}\ll 1$, the projectors become nearly identical and the left-right asymmetry is suppressed, signaling partial restoration of chiral symmetry.
Substituting these results into the expression for $C_{L}(\bm{k})$, we obtain
$C_{L}(\bm{k})=\nu_{k}\,\mathbb{I}_{2\times 2},\,\,\,\,\nu_{k}=\frac{1}{2}\left[\left(1-\frac{M_{q}}{E_{k}}\right)n_{+}+\left(1+\frac{M_{q}}{E_{k}}\right)(1-n_{-})\right].$ (13)
The coefficient $\nu_{k}$ plays the role of an effective occupation probability for the left-handed subspace. It incorporates both the thermal population ($n_{\pm}$) and the dynamical left-right mixing determined by $M_{q}/E_{k}$. From the positivity of $P_{L}\Lambda_{\pm}P_{L}$ and $0\leq n_{\pm}\leq 1$, it follows that $0\leq\nu_{k}\leq 1$, guaranteeing a valid probabilistic interpretation.
Physically, $\nu_{k}$ quantifies the degree of chirality coherence at momentum $k$:
  • For $\nu_{k}\simeq 1$ or 0, the state is dominantly left- or right-handed, indicating strong chiral polarization and negligible entanglement.
  • For $\nu_{k}\simeq\tfrac{1}{2}$, left- and right-handed components are equally mixed, reflecting maximal chirality entanglement and partial restoration of chiral symmetry.
In characteristic limiting cases at finite chemical potential $\mu_{q}$, we find
$M_{q}\to 0:$ $\,\,\,\,\nu_{k}=\tfrac{1}{2}\left(n_{+}+1-n_{-}\right),$ $\text{(chiral-symmetric limit)},$ (14)
$T\to 0:$ $\,\,\,\,\nu_{k}=\tfrac{1}{2}\left(1+\frac{M_{q}}{E_{k}}\right),$ $\text{(zero-temperature vacuum)},$ (15)
$T\to\infty:$ $\,\,\,\,\nu_{k}=\tfrac{1}{2},$ $\text{(maximal thermal mixing)}.$ (16)
Hence, $\nu_{k}$ evolves smoothly between the chirally broken and chirally symmetric regimes. In the vacuum, the finite constituent mass $M_{q}$ skews $\nu_{k}$ toward unity, signifying intense chiral polarization, while at high temperature or vanishing mass $\nu_{k}\to 1/2$, indicating the recovery of left-right symmetry and maximal entanglement.
Physically, $\nu_{k}$ quantifies the degree of left-right mixing in the quark spinor basis. When $M_{q}/E_{k}\ll 1$, the mass-induced coupling between left- and right-handed sectors weakens, yielding $\nu_{k}\simeq 1/2$ and indicating nearly maximal chirality entanglement. Conversely, for larger $M_{q}/E_{k}$, the occupation probability skews toward $nu_{k}\simeq 1$ or 0, signifying dominant chiral polarization and diminished entanglement. This transition from polarized to mixed chirality states encapsulates the dynamical mechanism of chiral symmetry restoration: as the dynamical mass decreases with increasing temperature or chemical potential, $\nu_{k}$ evolves continuously toward $1/2$, marking the onset of maximal quantum mixing between the two chiral sectors.
Fig. 2 illustrates the behavior of the effective occupation probability for the left-handed component, $\nu_{k}$ in Eq. (14), evaluated at $|\bm{k}|=1~\mathrm{GeV}$ for two representative constituent masses, $M_{q}=0.1~\mathrm{GeV}$ and $0.3~\mathrm{GeV}$. Panel (a) corresponds to the lighter quark mass, where $\nu_{k}$ approaches $1/2$ over a wide temperature range, while panel (b) represents a heavier constituent mass, exhibiting more substantial deviation from $1/2$.
V Von Neumann Entropy and Chirality Entanglement
The von Neumann entropy for the left-chiral subsystem quantifies the loss of quantum coherence between left- and right-handed components due to chiral mixing:
$S_{\chi}=-\mathrm{Tr}\left[C_{L}\ln C_{L}+(1-C_{L})\ln(1-C_{L})\right].$ (17)
This expression is formally identical to the von Neumann entropy in quantum information theory, but here the reduced density operator is replaced by the left-handed correlator $C_{L}$. The eigenvalues of $C_{L}$ encode how strongly the chiral subspaces are entangled. If $C_{L}$ is idempotent ($C_{L}^{2}=C_{L}$), the state is pure, and the entropy vanishes, whereas any deviation from idempotency indicates mixedness arising from chiral correlations.
For isotropic and spin-averaged matter, the correlator becomes diagonal, $C_{L}(\bm{k})=\nu_{k}\,\mathbb{I}_{2\times 2}$, simplifying the trace to
$\frac{S_{\chi}}{V}=2N_{c}N_{f}\int_{0}^{\Lambda}\frac{d^{3}k}{(2\pi)^{3}}\left[-\nu_{k}\ln\nu_{k}-(1-\nu_{k})\ln(1-\nu_{k})\right],$ (18)
where $\nu_{k}$ is given by Eq. (13). The integrand is positive and real for $0\leq\nu_{k}\leq 1$, representing the entropy contribution from each momentum mode $k$. The integration over all $\bm{k}$ combines these microscopic contributions into the macroscopic entanglement entropy density, $S_{\chi}/V$.
The quantity inside the integral can be identified as the single-mode chirality entropy,
$s_{\chi}(k)=-2\left[\nu_{k}\ln\nu_{k}+(1-\nu_{k})\ln(1-\nu_{k})\right].$ (19)
This expression is formally equivalent to the Shannon entropy of a binary probability distribution with occupation probabilities $\nu_{k}$ and $1-\nu_{k}$. It measures the degree of uncertainty or mixing between the left- and right-handed components for a given momentum mode:
  • $s_{\chi}(k)=0$ when $\nu_{k}=0$ or $1$, corresponding to a purely left- or right-handed state with perfect chirality coherence (no entanglement).
  • $s_{\chi}(k)$ reaches its maximum at $\nu_{k}=\tfrac{1}{2}$, indicating complete loss of chirality information and maximal entanglement.
As the constituent mass $M_{q}$ decreases, the mixing between chiral sectors becomes stronger, driving $\nu_{k}$ toward $\frac{1}{2}$ and enhancing $s_{\chi}(k)$. Consequently, the total entropy $S_{\chi}$ increases, reflecting the progressive loss of left-right coherence as the system approaches the chirally symmetric phase. In the chiral-symmetric limit $M_{q}\to 0$, $\nu_{k}\to\tfrac{1}{2}$, and the entropy attains its maximum value, corresponding to full quantum mixing between left and right components. In contrast, for the fully chirality-conserved vacuum ($M_{q}\neq 0$, $T=\mu=0$), the left and right subspaces remain orthogonal and unentangled, so that $\nu_{k}\to 1$ or 0 and $S_{\chi}\to 0$.
Hence, $S_{\chi}(T,\mu)$ serves as an information-theoretic probe for chiral symmetry restoration: Its rise with temperature or chemical potential signals the degradation of chiral coherence, while its vanishing at zero temperature and finite mass characterizes the chiral-broken, coherent vacuum. For reference, several limiting forms are listed below: At $T\!\to\!0$, $\nu_{k}=\theta(\mu_{q}-E_{k})$ and $s_{\chi}(k)=0$. At $T\!\to\!\infty$, $n_{F}\!\to\!\tfrac{1}{2}$ so $\nu_{k}=\tfrac{1}{2}$ and $s_{\chi}(k)=\ln 2$. In the chiral limit $M_{q}\!\to\!0$, $P_{L}\Lambda_{\pm}P_{L}=\tfrac{1}{2}P_{L}$, giving $\nu_{k}=\tfrac{1}{2}\,n_{F}(E_{k})$ and $s_{\chi}(k)=-\tfrac{1}{2}\!\big[n_{F}\ln n_{F}+(1-n_{F})\ln(1-n_{F})\big]$. These limits show that $s_{\chi}(k)$ vanishes for fully occupied or empty modes, approaches $\ln 2$ at high temperature, and varies smoothly near the chiral critical region.
VI Numerical Results and Discussions
            Refer to caption  (a)                         Refer to caption  (b)                         Refer to caption  (c)                         Refer to caption  (d)
Figure 3: Chirality entanglement entropy $S_{\chi}$ in Eq. (18), being normalized as $S_{\chi}/S^{(T,\mu)=0}_{\chi}$ as a function of $T$ and $\mu_{q}$ for $m_{q}=0$ (a) and $m_{q}=5.25$ MeV (b). Panels (c) and (d) show the corresponding 2-D slices at fixed $\mu_{q}$.
We now present the numerical evaluation of the chirality entanglement entropy density, $S_{\chi}/V$, as a function of temperature $T$ and quark chemical potential $\mu_{q}$ for both the chiral limit ($m_{q}=0$) and the explicitly broken case ($m_{q}=5.25~\mathrm{MeV}$). The results, displayed in Fig. 3(a-d), are computed from Eq. (18) using the self-consistent dynamical mass $M_{q}(T,\mu_{q})$ obtained from the NJL gap equation. Here, we show the normalized entropy as $S_{\chi}/S^{(T,\mu)=0}_{\chi}$. For instance, the absolute values are $S_{\chi}/V=(2.51,2.57)\times 10^{-2}~\mathrm{GeV}^{3}$ for $m_{q}=(0,5.25)$ MeV at $(T,\mu)=0$.
At all chemical potentials, $S_{\chi}/V$ increases monotonically with $T$, indicating a gradual loss of left-right coherence and an approach to a maximally mixed chiral state. Two microscopic mechanisms underlie this trend: (i) thermal broadening of the Fermi-Dirac distributions drives $n_{\pm}\to 1/2$, which pushes $\nu_{k}\to 1/2$ for all momentum modes and thus maximizes the single-mode entropy $s_{\chi}(k)$ and (ii) the dynamical mass $M_{q}(T,\mu_{q})$ decreases with increasing $T$, reducing chiral polarization and enhancing the coupling between left- and right-handed components. In the high-temperature regime, $M_{q}\to 0$ and $s_{\chi}(k)$ saturates, leading to a plateau in $S_{\chi}/V$ that corresponds to the chirally restored phase with maximal chirality mixing.
At fixed temperature, $S_{\chi}/V$ increases with $\mu_{q}$. A finite chemical potential raises the quark occupation probability below the Fermi surface ($n_{+}\simeq 1$ for $E_{k}<\mu_{q}$) and depletes antiquark states ($n_{-}\simeq 0$), broadening the momentum region where $\nu_{k}\approx 1/2$. This expansion of the near-maximal $s_{\chi}(k)$ domain yields a larger integrated entropy, resulting in $S_{\chi}(\mu_{q}=0)<S_{\chi}(\mu_{q}\neq 0)$ across the entire temperature range. This monotonic hierarchy parallels the thermodynamic enhancement of phase space and represents the partial restoration of chiral symmetry at finite density.
A comparison between Fig. 3(a) and Fig. 3(b), corresponding to $m_{q}=0$ and $m_{q}=5.25~\mathrm{MeV}$, respectively, shows that explicit symmetry breaking slightly suppresses $S_{\chi}/V$ at low and intermediate temperatures. A finite current mass keeps $M_{q}$ nonzero even near the crossover region, preserving partial left-right polarization and thereby reducing entanglement. However, as temperature rises, thermal fluctuations dominate the explicit breaking effects, and both cases converge to nearly identical entropy densities, consistent with the chiral-symmetric limit $M_{q}(T)\to 0$. Corresponding 2-D plots are also given in Figs. 3(c,d). In the chiral limit ($m_{q}\to 0$), the dynamical mass $M_{q}$ vanishes discontinuously across the second-order critical line, and the entropy $S_{\chi}$ exhibits a corresponding non-analytic jump in its temperature derivative. This reflects the singular change in the degree of left-right mixing at the phase boundary, where the system switches abruptly from a nearly pure chiral state to a maximally entangled one.
Overall, the monotonic rise of $S_{\chi}(T,\mu_{q})$ with both $T$ and $\mu_{q}$ demonstrates that the chirality entanglement entropy acts as an information-theoretic probe for chiral symmetry restoration. Unlike the conventional condensate $\langle\bar{q}q\rangle$, which measures the magnitude of symmetry breaking, $S_{\chi}$ quantifies the degree of quantum coherence loss between left- and right-handed sectors. In the chirally broken phase ($T\approx 0$, $M_{q}\neq 0$), the system is nearly pure with $\nu_{k}\approx 1$ or 0, and $S_{\chi}\simeq 0$. As the system passes through the pseudo-critical region, $\nu_{k}$ approaches $1/2$ near the Fermi surface, and $S_{\chi}$ rises sharply, marking the onset of a statistically mixed, chirally symmetric phase.
            Refer to caption  (a)                         Refer to caption  (b)
Figure 4: (a) Normalized $dS_{\chi}/dT$ (solid) and $|dM_{q}/dT|$ (dotted) as functions of temperature at $\mu=0$. The pseudo-critical temperatures extracted from the two quantities are $T_{\mathrm{pc}}^{S_{\chi}}=167~\mathrm{MeV}$ and $T_{\mathrm{pc}}^{M_{q}}=185~\mathrm{MeV}$, respectively, corresponding to a temperature difference of $\Delta T\simeq 18~\mathrm{MeV}$ between the two peaks. (b) Log-log scaling of the dynamical mass $M_{q}$ and the chirality entropy $S_{\chi}$ near the chiral critical temperature at $(m_{q},\mu_{q})=0$, i.e., $\ln\left[M_{q}(T)/\mathrm{GeV}\right]$ and $\ln\left[\Delta S_{\chi}(T)/\mathrm{GeV}^{3}\right]$ as functions of $\ln[\Delta T/\mathrm{GeV}]$, indicating the slopes $\beta_{M_{q}}\simeq 0.5$ and $\beta_{S_{\chi}}\simeq 1$. See the text for more details.
Fig. 4 compares the temperature derivatives of the chirality entanglement entropy and the dynamical quark mass, $dS_{\chi}/dT$ and $dM_{q}/dT$, at $\mu=0$ and $m_{q}=5.25$ MeV. Both show characteristic peaks near the pseudo-critical (pc) region, indicating rapid changes in chiral structure as the system approaches the restoration phase. The derivative $dS_{\chi}/dT$ exhibits a sharper peak at a slightly lower temperature than $|dM_{q}/dT|$, showing that the entanglement entropy is more sensitive to the onset of chiral decoherence. The difference between the two transition points reflects that quantum mixing between left- and right-handed sectors develops earlier than the complete melting of the dynamical mass. Physically, the rise and fall of $dS_{\chi}/dT$ represent the formation and saturation of chiral mixing as temperature increases. Once coherence between chiral sectors is fully lost, the entropy ceases to grow, marking the chirally restored regime. Hence, Fig. 4 demonstrates that $S_{\chi}$ acts as an entanglement susceptibility, providing an information-theoretic signal of the chiral transition complementary to the conventional order parameter $M_{q}(T)$.
Quantitatively, the peak positions of the two temperature derivatives in Fig. 4 differ by $\Delta T\simeq 18~\mathrm{MeV}$, corresponding to $T_{\mathrm{pc}}^{S_{\chi}}=167~\mathrm{MeV}$ and $T_{\mathrm{pc}}^{M_{q}}=185~\mathrm{MeV}$. This small but finite separation indicates that the onset of chiral quantum decoherence precedes the rapid melting of the dynamical quark mass as discussed previously. For $T>T_{\mathrm{pc}}$, the entanglement entropy $S_{\chi}(T)$ gradually saturates toward a normalized value $S_{\chi}/S_{\chi}^{\mathrm{max}}\approx 1$, signifying that the left- and right-handed sectors become maximally mixed in the chirally restored phase where further thermal excitation no longer changes the degree of quantum coherence.
From a quantum information viewpoint, thermal and density effects act as decohering agents that progressively erase phase correlations between chiral sectors. When $M_{q}\to 0$, this decoherence becomes complete, and each quark mode forms an equal superposition of left and right chirality, corresponding to maximal entanglement and maximal $S_{\chi}$. Thus, the numerical results confirm that $S_{\chi}(T,\mu_{q})$ continuously interpolates between a pure, chirally broken vacuum ($S_{\chi}\simeq 0$) and a fully mixed, chirally symmetric thermal state ($S_{\chi}\simeq S_{\chi}^{\mathrm{max}}$), establishing a direct and quantitative connection between chiral symmetry restoration and the growth of quantum entanglement in QCD matter. $S_{\chi}$ uniquely captures quantum coherence loss between chiral sectors, which is not accessed by conventional order parameters.
[…]
$M_{q}(T,\mu)=A_{M_{q}}\left(T^{M_{q}}_{c}-T\right)^{\beta_{M_{q}}},$ (20)
with the fitted parameters $A_{M_{q}}=1.38$, $\beta_{M_{q}}=0.52$, and $T^{M_{q}}_{c}=0.182~\mathrm{GeV}$. The value $\beta_{M}\simeq 1/2$ confirms the expected mean-field critical behavior in the chiral limit of the NJL model at $\mu_{q}=0$ [21].
To examine how the chirality entanglement entropy behaves in the same critical region, we introduce the analogous scaling form
$S_{\chi}(T,\mu)=S_{\chi}(T_{c},\mu)+A_{S_{\chi}}\left(T^{S_{\chi}}_{c}-T\right)^{\beta_{S_{\chi}}},$ (21)
where $S_{\chi}(T_{c},\mu)$ is the entropy at the critical point and $\beta_{S_{\chi}}$ denotes the entropy-scaling critical exponent. By fitting to the numerical results, we have the parameters $A_{S_{\chi}}=-0.143$, $\beta_{S_{\chi}}=1.01$, and $T^{S_{\chi}}_{c}=T^{M_{q}}_{c}$. For a reliable extraction of the critical exponents, the fit is restricted to the scaling window $0.172\,\mathrm{GeV}\lesssim T\lesssim T_{c}$, within which both $M_{q}$ and $S_{\chi}(T_{c})-S_{\chi}(T)$ exhibit clean linear behavior. All the relevant fitting parameters are listed in Table 2.
Quantity AA $\beta$ $T_{c}~[\mathrm{GeV}]$
$M_{q}$ $A_{M_{q}}=1.38$ $\beta_{M_{q}}=0.52$ $T^{M_{q}}_{c}=0.182$
$S_{\chi}$ $A_{S_{\chi}}=-0.143$ $\beta_{S_{\chi}}=1.01$ $T^{S_{\chi}}_{c}=0.182$
Table 2: Fitted scaling parameters for the dynamical mass $M_{q}$ and the chirality entanglement entropy $S_{\chi}$ in the chiral limit.
The dynamical mass $M_{q}$ acts as the chiral order parameter in the NJL model. In the chiral limit, the Ginzburg-Landau functional near the critical point is well approximated by $F[\phi]\propto a\,(T-T_{c})\phi^{2}+b\,\phi^{4}$, whose minimization gives the mean-field scaling $\phi\propto(T_{c}-T)^{1/2}$, resulting in $\beta_{M_{q}}=\frac{1}{2}$. This exponent reflects the universal critical behavior of the chiral order parameter.
The chirality entanglement entropy, however, is not an order parameter. Since it is defined through the thermodynamic derivative
$S_{\chi}=-\frac{\partial\Omega}{\partial T},$ (22)
and its scaling is governed by the temperature dependence of the free energy. Near the critical point, the dominant contribution behaves as $\Omega(T)\sim M_{q}^{2}(T)$, so that
$S_{\chi}(T_{c})-S_{\chi}(T)\propto(T_{c}-T)^{1},\,\,\,\,\beta_{S_{\chi}}\simeq 1.$ (23)
Thus the entropy exponent $\beta_{S_{\chi}}$ does not coincide with the order-parameter exponent $\beta_{M_{q}}$, leading to $\beta_{S_{\chi}}\approx 2\beta_{M_{q}}$ as in the mean-field analyses. While $S_{\chi}$ remains highly sensitive to the chiral transition, its scaling reflects thermodynamic derivatives rather than the universal order-parameter dynamics.
Fig. 4(b) displays the logarithmic scaling analysis of the dynamical mass $M_{q}(T)$ and the chirality entanglement entropy $S_{\chi}(T)$ in the chiral limit at $\mu_{q}=0$, i.e., $\ln\left[M_{q}(T)/\mathrm{GeV}\right]$ and $\ln\left[\Delta S_{\chi}(T)/\mathrm{GeV}^{3}\right]$ are given as functions of $\ln\left[\Delta T/\mathrm{GeV}\right]$, where $\Delta S_{\chi}(T)=S_{\chi}(T_{c})-S_{\chi}(T)$ and $\Delta T=T_{c}-T$. The clear difference between the two slopes in Fig. 4(b), therefore, highlights the distinct roles of the order parameter and its associated entanglement response: While $M_{q}$ encodes the universal mean-field criticality of SBCS, the chirality entropy quantifies the growth of quantum mixing between the chiral sectors. It thus inherits a modified, derivative-type scaling law.
VII Summary and Outlook
In this work, we have developed a consistent framework to quantify the quantum entanglement between left- and right-handed quark sectors within the finite-temperature and density Nambu–Jona-Lasinio (NJL) model. Starting from the lesser Green’s function $G^{<}(k)$ in the real-time formalism, we derived the left-handed reduced correlator $C_{L}(\bm{k})=P_{L}G^{<}(k)P_{L}$ and introduced the von Neumann chirality entropy, $S_{\chi}=-\mathrm{Tr}\!\left[C_{L}\ln C_{L}+(1-C_{L})\ln(1-C_{L})\right]$, as an information-theoretic measure of chiral coherence, this entropy characterizes the degree of mixing between the two chiral components, thereby providing a direct link between dynamical mass generation and the quantum information content of QCD matter.
The self-consistent solutions of the NJL gap equation reveal a QCD-like chiral phase structure: A second-order transition in the chiral limit and a smooth crossover for finite current quark mass. The dynamical quark mass $M_{q}(T,\mu_{q})$ decreases continuously with temperature and chemical potential, signaling progressive chiral symmetry restoration. Using this behavior, we computed the chirality entanglement entropy density $S_{\chi}/V$ and found that it increases monotonically with both $T$ and $\mu_{q}$, approaching a maximal value as $M_{q}\to 0$. Thus, $S_{\chi}$ mirrors the order-parameter behavior of $M_{q}$, i.e., it is nearly zero in the broken phase, where the system is chirally polarized, and maximal in the restored phase, where left- and right-handed components are thoroughly mixed.
A key aspect of this work is the scaling analysis of $S_{\chi}$ near the second-order critical line in the chiral limit. The dynamical mass exhibits the mean-field behavior $\beta_{M_{q}}\simeq 1/2$, being consistent with the standard Landau universality of the NJL model. In contrast, the entropy provides $\beta_{S_{\chi}}\simeq 1$, i.e., a distinctly larger exponent. This difference arises because $S_{\chi}$ is the temperature derivative of the free energy. At the same time, the critical part of the latter is dominated by the order-parameter amplitude squared, $\Omega_{\rm crit}\sim M_{q}^{2}\sim(T_{c}-T)^{1}$. Consequently, the entropy inherits the $M_{q}^{2}$ scaling rather than $M_{q}$, leading to an exponent approximately twice that of the order parameter. This confirms that $S_{\chi}$ is not an order parameter itself, but a thermodynamic entanglement response that nevertheless remains tightly correlated with the underlying critical dynamics.
The monotonic increase of $S_{\chi}(T,\mu_{q})$ therefore encodes the loss of left-right quantum coherence as the system evolves from a strongly correlated vacuum to a thermally randomized quark ensemble. In this sense, $S_{\chi}$ serves as an information-theoretic probe for chiral symmetry restoration, complementing the traditional condensate $\langle\bar{q}q\rangle$ by emphasizing the quantum entanglement structure of the quark sector. The framework developed here thus bridges thermal quantum field theory and quantum information theory, allowing symmetry restoration to be interpreted as the emergence of maximal quantum mixing in the chiral subspace.
The present study opens several directions for future work. Incorporating the Polyakov loop or explicit color interactions would enable a combined analysis of deconfinement and chirality entanglement. Extensions to PNJL or Dyson-Schwinger frameworks could clarify the interplay among confinement, dynamical mass, and information entropy. In confining models such as the PNJL framework, color-singlet constraints suppress single-quark excitations below the deconfinement temperature, reducing the accessible Hilbert-space volume and lowering $S_{\chi}$ in the confined phase. This suggests that confinement delays the onset of large chirality entanglement until color screening becomes effective near $T_{c}$. Future lattice-QCD simulations may directly evaluate fermionic correlation matrices or reduced chiral projectors to obtain $S_{\chi}$ nonperturbatively and thus test our entanglement-based interpretation of chiral symmetry restoration, since $C_{L}(\bm{k})$ is constructed from fermionic two-point functions and the same reduced correlator can in principle be evaluated on the lattice. Such developments would further establish the von Neumann chirality entropy as a universal, information-based diagnostic of phase transitions in strongly interacting quantum systems. Related works will appear elsewhere.
Acknowledgment
This work was supported by the National Research Foundation of Korea (NRF), funded by the Korean government (MSIT) (RS-2025-16065906).
Appendix
VII.1 Validity range of the equilibrium assumption
The equilibrium relations used in Eqs. (4)-(9),
$G^{<}(p)=in_{F}(p^{0})A(p),\,\,\,\,G^{>}(p)=-i[1-n_{F}(p^{0})]A(p),$ (24)
are valid when the system has reached local thermal equilibrium and its relaxation time $\tau_{\mathrm{rel}}$ is much longer than the microscopic interaction time. In this regime, quasiparticle excitations are well defined, and the spectral function $A(p)$ is sharply peaked at the mass shell,
$A(p)\simeq 2\pi\mathrm{sgn}(p^{0})\,\delta(p^{2}-M_{q}^{2}),$ (25)
so that the finite width of $A(p)$ can be neglected. This narrow-width (mean-field) limit justifies the use of the on-shell forms in Eqs. (6)-(9).
If a finite-width spectral function is employed,
$A(p)=\frac{-2\mathrm{Im}G^{R}(p)}{\pi},$ (26)
the definition of the lesser propagator and the reduced correlator, $C_{L}=P_{L}G^{<}P_{L}$, remains unchanged, but the occupation factor $\nu_{k}$ and the chirality entropy $S_{\chi}$ include an additional integration over the off-shell energy distribution:
$\nu_{k}=\int\frac{dp^{0}}{2\pi}n_{F}(p^{0})\mathrm{Tr}\left[P_{L}A(p)P_{L}\right].$ (27)
Consequently, $S_{\chi}$ measures the mixedness of the chiral subspace, including the effects of quasiparticle broadening, and reduces to the simple on-shell expression of Eq. (18) in the narrow-width limit. In this sense, the inclusion of a finite spectral width can be interpreted as an entanglement broadening effect, reflecting the enhanced mixing of chiral degrees of freedom due to thermal smearing and quasiparticle damping.
VII.2 Justification for using $C_{L}=P_{L}G^{<}P_{L}$ as a reduced density matrix
In equilibrium, the lesser Green’s function $G^{<}(k)$ describes the occupied part of the single-particle spectrum:
$C(\bm{k})=i\int\frac{dk^{0}}{2\pi}G^{<}(k),\qquad 0\leq C\leq 1,$ (28)
which defines the equal-time correlation matrix for a Gaussian (mean-field) fermionic state. Such a state is completely characterized by $C$, since higher-order correlations follow from Wick’s theorem.
Projecting onto the left-handed subspace via $P_{L}=(1-\gamma_{5})/2$, we obtain
$C_{L}=P_{L}CP_{L}=i\int\frac{dk^{0}}{2\pi}P_{L}G^{<}(k)P_{L},$ (29)
which represents the restriction of the full correlation matrix to the algebra generated by $\psi_{L}=P_{L}\psi$ and $\psi_{L}^{\dagger}$. The eigenvalues of $C_{L}$ satisfy $0\leq\nu_{k}\leq 1$ and play the role of effective occupation probabilities within the left-chiral sector.
For fermionic Gaussian states, the von Neumann entropy of the restricted (or “reduced”) subsystem is given by
$S_{\chi}=-\mathrm{Tr}\left[C_{L}\ln C_{L}+(P_{L}-C_{L})\ln(P_{L}-C_{L})\right],$ (30)
which coincides with the entropy of a true reduced density matrix. Hence, within the mean-field or quasiparticle approximation, $C_{L}$ serves as a legitimate surrogate for the reduced density operator of the left-handed sector. Its trace and spectrum are invariant under unitary basis changes, ensuring that $S_{\chi}$ is physically well-defined and basis independent.
This construction remains valid for any Gaussian equilibrium state with a positive spectral function $A(k)$ such that $G^{<}=in_{F}(k^{0})A(k)$. For interacting non-Gaussian cases, Eq. (30) provides the entropy of the Gaussian reference state sharing the same two-point correlator, serving as an upper bound to the exact subalgebra entropy.